A yo-yo is a solid cylindrical-shaped object wrapped by a string. if the yo-yo has a mass of 0.020kg and a radius of 10cm, find its speed when the string is unwound, moving downward at a distance of 1.0m. The unwinding of the string took place without slipping.
1
Expert's answer
2020-12-15T11:47:13-0500
mgh=0.5mv2+0.5Iω2I=0.5mr2mgh=0.5mv2+0.5(0.5mr2)ω2mgh=0.75mv2(9.8)(0.1)=0.75v2v=1.14msmgh=0.5mv^2+0.5I\omega^2\\I=0.5mr^2\\mgh=0.5mv^2+0.5(0.5mr^2)\omega^2\\mgh=0.75mv^2\\\\(9.8)(0.1)=0.75v^2\\v=1.14\frac{m}{s}